OpenStudy (anonymous):

Write the equation of the line that passes through (–3, 4) with a slope of 2 in point-slope form. Answer y + 3 = 2(x – 4) y + 4 = 2(x – 3) y – 3 = 2(x + 4) y – 4 = 2(x + 3)

OpenStudy (anonymous):

y-4=2(x+3)

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (saifoo.khan):

KGS ki fee kitne hai?

OpenStudy (anonymous):

tumhe ana hai kia !!???

OpenStudy (saifoo.khan):

That is not my answer.

hero (hero):

Go to bed Saifoo

OpenStudy (anonymous):

facebook par a !!

OpenStudy (saifoo.khan):

Good Morning Hero

OpenStudy (anonymous):

shut the .... up hero !!!

hero (hero):

I wasn't talking to Hash Brown

OpenStudy (saifoo.khan):

im online.

OpenStudy (anonymous):

u stalking on my gf

OpenStudy (saifoo.khan):

LOOL>

hero (hero):

You don't have a gf

OpenStudy (saifoo.khan):

Hashir?

OpenStudy (anonymous):

i have 4

OpenStudy (anonymous):

now go to hell hero

OpenStudy (anonymous):

hero=zero

hero (hero):

That's grounds for suspension

OpenStudy (anonymous):

and i dont have online gfs like u i have them i real u ......

OpenStudy (saifoo.khan):

hashir bta na bhai/

hero (hero):

I wasn't even talking about online girls

OpenStudy (anonymous):

facebook par !! main online hun

hero (hero):

That's all you

OpenStudy (saifoo.khan):

yar sirf fee btado.. aur kuch nahi!

OpenStudy (anonymous):

hero=zero u seem like a gadha

OpenStudy (anonymous):

gadha= suerstar genius

hero (hero):

My name is Hero, not hero = zero

hero (hero):

Saifoo must be way smarter than you. He knows not to participate in this non-sense